OpenStudy (anonymous):

Which is the smallest number? eleven twenty-fifths nine-twentieths three-fifths zero point five two

OpenStudy (anonymous):

\[11\frac{ 2 }{ 5 }\] \[\frac{ 9 }{ 20 }\] \[\frac{ 3 }{ 5 }\] \[0.\]

OpenStudy (anonymous):

\[0.52\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

***11/25 = 0.44 9/20 = 0.45 3/5 = 0.6 0.52